Graph the following system of equations and find any solutions. Check the answers. Identify the system as consistent or incon is this consistent or inconsistent? if it is consistent, state wether the equations are dependent or independent.
Answer #1

2- 3y = -18 -119 -3x+y= 6 - V0) 726) » X is [10,6) : The solution o apowion, it is consistendo * Since there is only one solu....